### v1

# Guillaume Balezo

Abstract
Atypical mitotic figures (AMFs) are markers of abnormal cell division associated with poor prognosis, yet their detection remains difficult due to low prevalence, subtle morphology, and inter-observer variability. The MIDOG 2025 challenge introduces a benchmark for AMF classification across multiple domains. In this work, we evaluate the recently published DINOv3-H+ vision transformer, pretrained on natural images, which we fine-tuned using low-rank adaptation (LoRA, 650k trainable parameters) and extensive augmentation. Despite the domain gap, DINOv3 transfers effectively to histopathology, achieving a balanced accuracy of 0.8871 on the preliminary test set. These results highlight the robustness of DINOv3 pretraining and show that, when combined with parameter-efficient fine-tuning, it provides a strong baseline for atypical mitosis classification in MIDOG 2025.
keywords:
histopathology | foundation models | atypical mitotic figures
{corrauthor}
guillaume.balezo@minesparis.psl.eu
Introduction
Mitotic activity is a central indicator of tumor proliferation and prognosis. Beyond simple counts, the distinction between normal mitotic figures (NMFs) and atypical mitotic figures (AMFs) is of particular interest, as AMFs reflect abnormal cell division processes and correlate with poor clinical outcome. However, their identification is challenging due to low prevalence, subtle morphological differences, and low inter-rater agreement even among trained pathologists. Automated image analysis methods therefore have the potential to improve reproducibility and reduce observer bias in this task.
The Mitosis Domain Generalization Challenge 2025 (MIDOG25) extends the scope of previous editions (MIDOG 2021 (1) and MIDOG 2022 (2)) with the goal of advancing AI-assisted cancer diagnosis. The Task 2 introduces a dedicated benchmark for AMF classification, where participants are asked to classify cropped cell patches (128×128 pixels) into NMF or AMF across multiple tumor types, species, scanners, and laboratories. The dataset comprises more than 12,000 annotated mitotic figures, with AMFs accounting for only $\sim$20% of cases, and the evaluation metric of the challenge is the balanced accuracy to mitigate the strong class imbalance. Similar to the earlier MIDOG challenges, this benchmark addresses the crucial problem of robustness and generalization across domains, now extended to the clinically relevant task of atypical mitosis classification.
In this work, we tackle the Task 2 by applying low-rank adaptation (LoRA) (3) to fine-tune the recently published DINOv3-H+ vision transformer (ViT) (4), a model pretrained on natural images using DINOv3 self-supervised (SSL) method. To enhance robustness across diverse domains and compensate for the limited number of atypical figures, we combine this strategy with extensive data augmentation. Our approach aims to test and leverage the representational power of this new self-supervised pretraining, while ensuring efficient adaptation to the heterogeneous challenge dataset.
Material and Methods
0.1 Dataset
The MIDOG 2025 atypical mitosis training set is derived from 454 histopathology images spanning nine domains defined by different tumor types, species, scanners, and laboratories. Each mitotic figure was subtyped as normal or atypical by three expert pathologists in a blinded majority-vote setting.
In addition to the official MIDOG 2025 atypical training set, we incorporated three external resources. The AMi-Br (5) dataset provides mitotic figures from MIDOG 2021 (1) and TUPAC16 (6); to avoid overlap, we only used the TUPAC16 cohort. The AtNorM-Br (7) dataset contains mitotic figures from the TCGA (8) breast cancer cohort, annotated by an expert pathologist. Finally, the OMG-Octo dataset (9) was created by screening large histopathology data with a model pretrained on AMi-Br and MIDOG25, followed by expert review of candidate mitoses.
After removing duplicate images, our training set comprised 11,939 mitotic figures from MIDOG 2025 (10,191 normal, 1,748 atypical), 1,999 mitotic figures from AMi-Br (1,571 normal, 428 atypical), 711 from AtNorM-Br (587 normal, 124 atypical), and 1,752 from OMG-Octo (378 normal, 1,374 atypical), resulting in a total of 16,398 figures (12,724 normal and 3,674 atypical). All datasets were provided as 128×128 pixel crops centered on the mitotic figure, except OMG-Octo, which was originally 64×64 pixels and resized to 128×128 for training, corresponding to a resolution of $0.25\,\mu\text{m/pixel}$.
The preliminary test set provided for the Task 2 consisted of mitotic figure crops from four tumor types not included in the final test data. It was made available on the challenge platform two weeks prior to submission for debugging purposes. The final test set consisted of 120 cases covering 12 distinct tumor types from both human and veterinary pathology, with 10 cases per tumor type. This set spans multiple laboratories and scanning systems and was used for the official evaluation. Performance was assessed using balanced accuracy, computed over all patches of the test set.
0.2 Network Training
We trained our model on $\times128$ pixel image crops, matching the challenge’s original patch size. Our model is a DINOv3-H+ vision transformer pretrained on the LVD-168M natural image dataset, which we fine-tuned for the Task 2 with low-rank adaptation (LoRA; $rank=4$, $\alpha_{\text{LoRA}}=8.0$, dropout 0.05, applied only to the query and value projections in the attention layers), resulting in only about 650k trainable parameters. A linear classification head with 0.5 dropout was added to produce logits from the class token. Training was run with a batch size of 16 and mixed precision (FP16). We optimized with AdamW (learning rate $1\times 10^{-4}$, weight decay 0.1, $\epsilon=1\times 10^{-7}$), using a cosine schedule with linear warmup during the first 10% of training (from $8.47\times 10^{-7}$ to the base rate). Gradient norms were clipped at 1.0 for stability.
To address the class imbalance ($\sim$20% atypical), we trained with Focal Loss (10) ($\alpha=0.25$, $\gamma=2$). Extensive online augmentations were applied, including color jitter, JPEG compression, stain augmentation (multi-Macenko (11, 12) with random stain domain references), defocus blur, affine transforms, D4 symmetry, coarse dropout (up to two random boxes), and a custom black-border augmentation to mimic zero-padded regions in the training data. Inputs were normalized with ImageNet statistics, consistent with DINOv3 pretraining. The final submitted model was trained for 60 epochs using all available datasets (AMi-Br TUPAC16, MIDOG25, AtNorM-Br, and OMG-Octo). At inference, we employed test-time augmentation by averaging logits across four rotated views to improve robustness.
The multi-Macenko augmentation used 10 stain references per domain, extracted from the training set as well as MITOS CMC (13), MITOS CCMCT (14), and TCGA COAD/BLCA cases. During training, a domain and number of references were sampled at random to mimic diverse staining conditions.
Evaluation and Results
We evaluated our method with 4-fold cross-validation on the complete training data, holding out the AMi-Br TUPAC16 subset as an external test set. In each fold, the model was evaluated on the validation fold, as well as on the AMi-Br TUPAC16 test set and the preliminary test set. We report mean and standard deviation across folds for balanced accuracy (BA).
Table 1: Performance of our method on 4-fold cross-validation (mean ± std), the external AMi-Br (TUPAC) test set, and the preliminary test set.
Split BA
Cross-validation 0.927 $\pm$ 0.002
AMi-Br TUPAC test 0.842 $\pm 0.004$
Preliminary Test Set 0.887
We also explored continuing the self-supervised training of the DINOv3-L (LVD) model on mitosis-like images obtained from an object detector trained on Task 1. Candidate patches were collected from TCGA BLCA, COAD, MITOS CMC, MITOS CCMCT, and the challenge training data, resulting in a dataset of $\sim$260k crops (128×128). Using the official DINOv3 pipeline, we fine-tuned LoRA parameters starting from ImageNet-pretrained weights. Due to limited compute (4 GPUs, batch size 4 vs. the original 2048), training was constrained. Linear probing on the training set showed that this additional pretraining yielded a relative improvement of approximately 10% in linear balanced accuracy, increasing from 53.44 % with ImageNet initialization to 63.11 %. This suggests that large-scale SSL on mitosis images could be beneficial. However, due to time and budget constraints, we did not attempt full LoRA fine-tuning of this model, leaving it as a promising direction for future work.
Discussion
In this work, we showed that DINOv3-H+ with LoRA fine-tuning provides a strong baseline for atypical mitosis classification in MIDOG 2025, while requiring training of only $\sim$650k parameters. The robust pretraining of DINOv3 on natural images appears to transfer well to histopathology, even though the domain shift is substantial. On the preliminary test set, our model achieved a balanced accuracy of 0.8871. These results suggest that parameter-efficient adaptation combined with extensive augmentation can achieve competitive performance under limited data and severe class imbalance.
Future work could extend this study with broader validation and exploration of alternative DINOv3 variants (e.g., Large, 7B) and large-scale self-supervised pretraining directly on histopathology data. Such efforts may further improve generalization across domains and strengthen the applicability of foundation models for mitosis subtyping.